OpenStudy (anonymous):

Let f(x) = x – 2 and g(x) = x2 – 7x – 9. Find f(g(–1)).

OpenStudy (asnaseer):

ok, first evaluate g(-1)

OpenStudy (asnaseer):

by using x=-1 in the expression for g(x)

OpenStudy (asnaseer):

what do you get?

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

@dpaInc yes please!

OpenStudy (anonymous):

\(\large f(x) = x-2 \), \(\large g(x)=x^2-7x-9 \) you want \(\large f(g(-1)) \).. now, \(\large g(-1)=(-1)^2-7(-1)-9=1+7-9=-1 \) so, \(\large f(g(-1))=f(-1)=-1-2=-3 \)

OpenStudy (anonymous):

it's ok if you have questions... ask 'em...:)